# s1: Simple test-time scaling

### v1

@@ B.2 Dataset composition for full 59K questions
OlympiadBench (He et al., 2024b) Math and Physics olympiad questions 896 3.9K
@@
LiveCodeBench (Jain et al., 2024b) Code problems from coding websites (LeetCode, AtCoder, and CodeForces) 151 3.5K
@@
10:        Select correct MATH solutions with long chains
@@ Appendix C Training details
We take a model that has already been pretrained and instruction tuned and further finetune it for reasoning. Specifically, we use Qwen2.5-32B-Instruct (Qwen et al., 2024), which generally matches or outperforms the larger Qwen2.5-72B-Instruct on math tasks (Qwen et al., 2024). We use token delimiters to separate the thinking stage from the answering stage. We enclose the thinking stage with <|im_start|>think and <|im_start|>answer; both preceded and followed by a newline. Samples from our dataset are in §C.2. We use basic fine-tuning hyperparameters: we train for 5 epochs with a batch size of 16 for a total of 315 gradient steps. We train in bfloat16 precision with a learning rate of $51e-$ 5 warmed up linearly for 5% (16 steps) and then decayed to 0 over the rest of training (299 steps) following a cosine schedule. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9,\beta_{2}=0.95$ and weight decay of $41e-$ 4. We do not compute loss on questions, only on reasoning traces and solutions. We ensure the sequence length is large enough to avoid cutting off any samples; a setting we ablate in §C.1. The training takes just 26 minutes on 16 NVIDIA H100 GPUs.
@@ C.1 Training Ablations: Sequence length
Besides our scaling ablations in §5.2, the main training hyperparameter we ablate is the sequence length used during training. We find that a shorter training sequence length leads to longer reasoning traces at test time. This is because when training with a shorter sequence length the answer section of the training sample is more commonly cut off. Inversely, when the training sequence length is longer, more samples appear in their entirety with the section where the model answers. Thus the model receives more gradient updates where it learns to generate an answer following its chain. This in turn leads to a higher log probability of the answer section at any point during the generation and thus shorter reasoning traces at test time. Performance-wise, we find that there is no major difference when training with different sequence lengths, thus we opt for the longest training sequence length as it makes inference more efficient by leading to shorter reasoning traces.
@@ D.1 Sequential scaling ablations
Prompt appended to the question after two newlines AIME24 MATH GPQA

### v2

@@ B.2 Dataset composition for full 59K questions
OlympiadBench (He et al., 2024) Math and Physics olympiad questions 896 3.9K
@@
LiveCodeBench (Jain et al., 2024) Code problems from coding websites (LeetCode, AtCoder, and CodeForces) 151 3.5K
@@
10:        Select correct MATH500 solutions with long chains
@@ Appendix C Training details
We take a model that has already been pretrained and instruction tuned and further finetune it for reasoning. Specifically, we use Qwen2.5-32B-Instruct (Qwen et al., 2024), which on math tasks generally matches or outperforms the larger Qwen2.5-72B-Instruct (Qwen et al., 2024) or other open models (Dubey et al., 2024; Groeneveld et al., 2024; Muennighoff et al., 2024). We use token delimiters to separate the thinking stage from the answering stage. We enclose the thinking stage with <|im_start|>think and <|im_start|>answer; both preceded and followed by a newline. Samples from our dataset are in §C.2. We use basic fine-tuning hyperparameters: we train for 5 epochs with a batch size of 16 for a total of 315 gradient steps. We train in bfloat16 precision with a learning rate of $51e-$ 5 warmed up linearly for 5% (16 steps) and then decayed to 0 over the rest of training (299 steps) following a cosine schedule. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9,\beta_{2}=0.95$ and weight decay of $41e-$ 4. We do not compute loss on questions, only on reasoning traces and solutions. We ensure the sequence length is large enough to avoid cutting off any samples; a setting we ablate in §C.1. The training takes just 26 minutes on 16 NVIDIA H100 GPUs.
@@ C.1 Training Ablations: Sequence length
Besides our scaling ablations in §5.2, the main training hyperparameter we ablate is the sequence length used during training. We find that a shorter training sequence length leads to longer reasoning traces at test time. This is because when training with a shorter sequence length the answer section of the training sample is more commonly cut off. Inversely, when the training sequence length is longer, more samples appear in their entirety with the section where the model answers. Thus the model receives more gradient updates where it learns to generate an answer following its chain. This in turn leads to a higher log probability of the answer section at any point during the generation and thus shorter reasoning traces at test time. Performance-wise, we also find that the model trained with a longer sequence length performs better. Thus we opt for the longest training sequence length as it leads to better performance and makes inference more efficient by leading to shorter reasoning traces.
@@ D.1 Sequential scaling ablations
Prompt appended to the question after two newlines AIME24 MATH500 GPQA